OpenStudy (samara89):

A non-scale model of the earth has a diameter of 18 in. Within an outer clay layer, there is a core made of a ball with a diameter of 3 in. What is the volume of the clay? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (samara89):

@greatlife44

OpenStudy (samara89):

I think the answer is 900.13

OpenStudy (greatlife44):

\[(4/3)\pi(r^3)_{ball} = (4/3)(\pi)(1.5)^{3} = 14.3 cm^{3}_{volume~ball}\] \[\frac{ 4 }{ 3 }*(9)^{3}*\pi = 3052~inches^{3}\] \[Volume~model-Volume~ball = 3052-14.3 = 3038inches^{3}\]

rebeccaxhawaii (rebeccaxhawaii):

do you understand?